# Continuity and equivariant dimension

### v1

@@ 1 Introduction
In the general noncommutative setting, key elements of the above discussion translate into conditions on freeness or equivariant dimension. Precursors to the study of equivariant dimension appeared in [55, §3], which motivated some later problems in noncommutative Borsuk-Ulam theory. Taghavi’s questions primarily concerned counting the number of odd self-adjoint elements required to produce an invertible square-sum. It should be noted that this is not the only interpretation of noncommutative Borsuk-Ulam theory; references [60, 3, 39, 11] consider an approach focused on (non-)existence of equivariant morphisms. The full formulation of equivariant dimension, including a much more general context, appeared in [19]. The local-triviality dimension of an action is defined in [19, Definition 3.1], and it is a noncommutative analogue of the Schwarz genus. Moreover, it is related to the Rokhlin dimension of [23, Definition 1.1], as explored in [19, §5.2]. We recall the definitions of the local-triviality dimension in Section 3.
@@ 2 Preliminaries
For what follows, $\mathrm{Prim}(A)$ denotes the primitive spectrum [42, §4.1.2] of a $C^{*}$-algebra $A$: the space of primitive ideals, equipped with the Jacobson (or hull-kernel) topology [42, §4.1.4]. Further, $Z(A)$ denotes the center of $A$, and $M(A)$ denotes the multiplier algebra of $A$. Following [45, §3], say, we write
@@
More generally, [35, §5] shows that the cyclic group $K_{1}(C({\mathbb{S}}^{2n-1}_{\theta}))\cong{\mathbb{Z}}$ admits a generator of matrix dimension $2^{n-1}$, which we will denote $Z_{\theta}$, such that each entry is a ***-monomial. For each of the ${\mathbb{Z}}/k$ rotation actions defined above, the eigenspace of each entry is such that there exist unitary matrices $A,B\in M_{2^{n-1}}({\mathbb{C}})$, independent of $\theta$, such that $R(Z_{\theta})=AZ_{\theta}B$ (where $R$ is applied entrywise) and $A^{k}=B^{k}=I$. The antipodal action ($k=2$) is particularly simple; in this case each entry of $Z_{\theta}$ is odd, so $R(Z_{\theta})=-Z_{\theta}$. See [39, §3] for a slightly modified construction that chooses the monomial coefficients continuously.
@@
The $\theta$-spheres are related to a similar family of $C^{*}$-algebras called quantum (or noncommutative) tori ([20, §12.2] or [47, §1]). If $\theta$ is an antisymmetric $n\times n$ real matrix, the $C^{*}$-algebra $A^{n}_{\theta}$, or $C(\mathbb{T}^{n}_{\theta})$, is defined by generators $u_{1},\ldots,u_{n}$ and relations
@@
All $\theta$-spheres and noncommutative tori are known to be Rieffel deformations in the sense of [49, §10.2]. While it is not clear if the $\theta$-spheres or noncommutative tori form a continuous field over the set of antisymmetric matrices, [49, Theorem 8.13] implies that the field is separately continuous in each entry. In particular, the distinction between those two cases does not matter for $3$-spheres or $2$-tori. As we will discuss continuity of fields in more detail, we recall some basic definitions.
@@
$C(X)$-algebras are one way to formalize the notion of a field of $C^{*}$-algebras [46, Definition 1.1 and Propositions 1.2 and 1.3], while [37, Theorem 2.3] equates the notion to that of an upper semicontinuous $C^{*}$-algebra bundle, generalizing [18, Definition VII.8.2].
@@ 3 Local-Triviality Dimension and Fields
The local-triviality dimension $\mathrm{dim}^{{\mathbb{G}}}_{\mathrm{LT}}(A)$ is the infinimum of the set of $n\in\mathbb{N}$ such that there exist ${\mathbb{G}}$-equivariant ***-homomorphisms $\rho_{0},\ldots,\rho_{n}:C_{0}((0,1])\otimes C({\mathbb{G}})\to A$ with $\sum\limits_{j=0}^{n}\rho_{j}(t\otimes 1)=1$. $\blacklozenge$
@@
The weak local-triviality dimension $\mathrm{dim}^{{\mathbb{G}}}_{\mathrm{WLT}}(A)$ is the infinimum of the set of $n\in\mathbb{N}$ such that there exist ${\mathbb{G}}$-equivariant ***-homomorphisms $\rho_{0},\ldots,\rho_{n}:C_{0}((0,1])\otimes C({\mathbb{G}})\to A$ with $\sum\limits_{j=0}^{n}\rho_{j}(t\otimes 1)$ invertible. $\blacklozenge$
@@
The strong local-triviality dimension is the infimum of the set of $n\in\mathbb{N}$ such that there exists a ${\mathbb{G}}$-equivariant unital ***-homomorphism $\rho:C(E_{n}^{\Delta}{\mathbb{G}})\to A$. $\blacklozenge$
@@
Multiple characterizations of freeness are also found in [4, Theorem 0.4]. Moreover, finiteness of any one of the local-triviality dimensions implies freeness of the coaction by [19, Theorem 3.8]. We will primarily be interested in the case that $A$ is noncommutative, but ${\mathbb{G}}$ is an ordinary group (in particular, a compact abelian group). In this case, there is an equivalent condition to freeness found in [48, Definition 1.6] or [43, Definition 7.1.4]. If ${\mathbb{G}}$ is compact abelian and $\alpha$ is an action of ${\mathbb{G}}$ on $A$, then for each character $\gamma\in\widehat{{\mathbb{G}}}$, there is an associated eigenspace (or spectral subspace)
@@
Immediate, as there exist ${\mathbb{G}}$-equivariant morphisms $A\to A_{x}$ for all $x\in X$. normal-$\blacksquare$
@@
is thus a self-map of the sphere that has no fixed points and sends no point to its antipode, and no such map exists by [57, Corollary 1.24]. $\blacklozenge$
@@
Equip ${\mathbb{P}}^{1}$ with its complex analytic structure or, equivalently for the purpose of manipulating holomorphic bundles, its complex projective algebraic structure (see [52, §12]). The line bundle
@@
Having to work with such trivializations extensively, we need some language to express the extent to which they exist. To that end, the following notions specialize the (numerical) ${\mathbb{G}}$-index ([61, Definition 3.13], [30, Definition 6.2.3]) of a ${\mathbb{G}}$-action on a space $X$, which is typically compact. That invariant can be recast via [13, Theorem 3.4] (where the term is co-index instead) as the smallest $n$ for which $X/{\mathbb{G}}$ admits an open cover by $U_{0},\ldots,U_{n}$, above which the action induces trivial principal ${\mathbb{G}}$-bundles.
@@
For a ${\mathbb{Z}}/k$-action on a matrix bundle ${\mathcal{M}}$ over $X$, the corresponding index $\operatorname{ix}({\mathcal{M}},\triangleright)=\operatorname{ix}(\triangleright)$ is the smallest $m$ for which there is an open cover $X=\bigcup_{0}^{m}U_{i}$ with ${\mathcal{M}}$ trivial, $\triangleright$ inner over each $U_{i}$ induced by some unitary $s_{i}\in\Gamma({\mathcal{M}}|_{U_{i}})$, $s_{i}^{k}=1$, with all eigenspaces of $s_{i}$ again trivial as bundles over $U_{i}$. $\blacklozenge$
@@
Given a ${\mathbb{Z}}/k$-action normal-$\triangleright$ on a matrix bundle ${\mathcal{M}}\to X$ over a compact Hausdorff space, we have
$\left(\operatorname{ix}({\mathcal{M}},\triangleright)+1\right)(\operatorname{\mathrm{qt}}+1)-1\,\,\geq\,\,\mathrm{dim}^{{\mathbb{Z}}/k}_{\mathrm{WLT}}(\Gamma({\mathcal{M}}))\,\,\geq\,\,\operatorname{\mathrm{qt}}.$ (3-4)
@@
[12, Theorem 4.1] effectively proves the singleton-$X$ version, when ${\mathcal{M}}$ is just a matrix algebra, $\operatorname{ix}:=\operatorname{ix}({\mathcal{M}},\triangleright)=0$, and 3-4 is a single equality. Although that discussion focuses on gauge actions on matrix algebras, the arguments apply to arbitrary conjugation actions by unitaries $v\in M_{n}$, $v^{k}=1$. This same remark, given the equivariant maps
$\Gamma({\mathcal{M}})=:A\xrightarrow{\quad}A_{x}\cong M_{n_{x}},$
also proves the lower bound $\mathrm{dim}^{{\mathbb{Z}}/k}_{\mathrm{WLT}}(A)\geq\operatorname{ix}$ in 3-4.
As for the upper bound on $\mathrm{dim}^{{\mathbb{Z}}/k}_{\mathrm{WLT}}(A)$, consider an open cover $X=\bigcup_{i}U_{i}$ trivializing the matrix bundle ${\mathcal{M}}$ and rendering the action inner, by unitary sections $v_{i}\in\Gamma({\mathcal{M}}|_{U_{i}})$ with $v_{i}^{k}=1$, and trivial $v_{i}$-eigenbundles
${\mathcal{E}}_{i,j}:=\ker(\zeta^{j}-v_{i})\quad 0\leq i\leq\operatorname{ix}\quad j\in{\mathbb{Z}}/k,$
uniquely determined (for each $i$) up to cycling through the $j$ indices. We also fix a refinement $(V_{i})$ of the cover $(U_{i})$, with
@@
Along each $U_{i}$ there are $\operatorname{\mathrm{qt}}+1$ operators $a_{i,s}\in A|_{U_{i}}=\Gamma({\mathcal{M}}|_{U_{i}})$, $0\leq s\leq\operatorname{\mathrm{qt}}$, each mapping
${\mathcal{E}}_{i,j}|_{U_{i}}\xrightarrow{\quad a_{i,s}\quad}{\mathcal{E}}_{i,j+1}|_{U_{i}},\quad j\in{\mathbb{Z}}/k$
partially isometrically with maximal possible rank $\min_{j}\operatorname{\mathrm{rk}}{\mathcal{E}}_{i,j}$, with $a_{i,s}^{k}$ a projection, and with $\sum_{s}a_{i,s}^{*}a_{i,s}$ invertible on each ${\mathcal{E}}_{i,j}|_{U_{i}}$. It remains to rescale $a_{i,s}$ into $\phi_{i}a_{i,s}\in A$ for continuous functions $\phi_{i}$ supported in $U_{i}$ with $\phi_{i}|_{V_{i}}\equiv 1$. normal-$\blacksquare$
@@
For a ${\mathbb{Z}}/k$-action normal-$\triangleright$ on a matrix bundle ${\mathcal{M}}\to X$ over a compact Hausdorff space, the following conditions are equivalent.
@@
a normal-$\Longrightarrow$ b follows from the upper bound 3-4 on $\mathrm{dim}^{{\mathbb{Z}}/k}_{\mathrm{WLT}}(\Gamma({\mathcal{M}}))$ while b normal-$\Longrightarrow$ c generally [19, Theorem 3.8]. As for c normal-$\Longrightarrow$ a, freeness means [33, Theorem 8.1.7] that the ${\mathbb{Z}}/k$-grading on $A:=\Gamma({\mathcal{M}})$ induced by the action is strong:
@@
then the orthogonal projection onto the latter (non-zero) space annihilates $M_{n,\zeta}M_{n,\zeta^{-1}}$. normal-$\blacksquare$
@@
For a ${\mathbb{Z}}/k$-action normal-$\triangleright$ on a matrix bundle ${\mathcal{M}}\to X$ over a compact Hausdorff space, the following conditions are equivalent.
@@
    $\mathrm{dim}^{{\mathbb{Z}}/k}_{\mathrm{LT}}(\Gamma({\mathcal{M}}))\leq\operatorname{ix}({\mathcal{M}},\triangleright)$;
@@
    $\mathrm{dim}^{{\mathbb{Z}}/k}_{\mathrm{LT}}(\Gamma({\mathcal{M}}))<\infty$;
@@
The implication a normal-$\Rightarrow$ b is formal and b normal-$\Rightarrow$ c follows from (the proof of) [12, Theorem 4.4], which provides the $X=\{*\}$ version; we thus turn to c normal-$\Rightarrow$ a.
As in the proof of Theorem 3.13, consider an open cover $X=\bigcup_{i=0}^{\operatorname{ix}}U_{i}$ trivializing ${\mathcal{M}}$, generators $v_{i}$ for the ${\mathbb{Z}}/k$-action restricted to $U_{i}$, and the eigenspaces ${\mathcal{E}}_{i,j}$, $j\in{\mathbb{Z}}/k$ of $v_{i}$. We can as in that earlier proof, the only difference being that now we can select a single unitary $a_{i}\in\Gamma({\mathcal{M}}|_{U_{i}})$ for each $i$, rescaling it to $\phi_{i}^{1/2}a_{i}$ for a partition of unity [34, Theorem 36.1] $(\phi_{i})$ subordinate to the cover $(U_{i})_{i}$. normal-$\blacksquare$
@@
[…]
@@
Because (by oddness) each eigenspace of $a_{i}$ is orthogonal to each eigenspace of $u$, 3-5 coincides with the canonical surjection ${\mathbb{S}}^{2}\to{\mathbb{R}}{\mathbb{P}}^{2}$ nowhere. A continuous lift ${\mathbb{S}}^{2}\to{\mathbb{S}}^{2}$ thereof (which does exist by [34, Lemma 79.1], as ${\mathbb{S}}^{2}$ is simply-connected) will thus have no coincidences with either the identity or the antipodal map, once more violating [57, Corollary 1.24]. normal-$\blacksquare$
@@
by projectivity, as in the proof of [12, Proposition 5.4], via [44, Proposition 2.10]. Define $a:=\sum\limits_{i=0}^{n}\overline{\rho}_{i}(t\otimes 1)$,. Note that $\pi_{x}(a)$ is invertible, so we may choose an element $b\in A$ such that $\pi_{x}(b)=(\pi_{x}(a))^{-1}$. That is, both $\|ab-1\|_{x}=0$ and $\|ba-1\|_{x}=0$. By [46, Proposition 1.2], for each $c\in A$ the mapping

### v2

@@ 1 Introduction
In the general noncommutative setting, key elements of the above discussion translate into conditions on freeness or equivariant dimension. Precursors to the study of equivariant dimension appeared in [56, §3], which motivated some later problems in noncommutative Borsuk-Ulam theory. Taghavi’s questions primarily concerned counting the number of odd self-adjoint elements required to produce an invertible square-sum. It should be noted that this is not the only interpretation of noncommutative Borsuk-Ulam theory; references [61, 3, 39, 41, 10, 11] consider an approach focused on (non-)existence of equivariant morphisms. The full formulation of equivariant dimension, including a much more general context, appeared in [19]. The local-triviality dimension of an action is defined in [19, Definition 3.1], and it is a noncommutative analogue of the Schwarz genus. Moreover, it is related to the Rokhlin dimension of [23, Definition 1.1], as explored in [19, §5.2]. We recall the definitions of the local-triviality dimension in Section 3.
@@ 2 Preliminaries
For what follows, $\mathrm{Prim}(A)$ denotes the primitive spectrum [43, §4.1.2] of a $C^{*}$-algebra $A$: the space of primitive ideals, equipped with the Jacobson (or hull-kernel) topology [43, §4.1.4]. Further, $Z(A)$ denotes the center of $A$, and $M(A)$ denotes the multiplier algebra of $A$. Following [46, §3], say, we write
@@
More generally, [35, §5] shows that the cyclic group $K_{1}(C({\mathbb{S}}^{2n-1}_{\theta}))\cong{\mathbb{Z}}$ admits a generator of matrix dimension $2^{n-1}$, which we will denote $Z_{\theta}$, such that each entry is a *-monomial. For each of the ${\mathbb{Z}}/k$ rotation actions defined above, the eigenspace of each entry is such that there exist unitary matrices $A,B\in M_{2^{n-1}}({\mathbb{C}})$, independent of $\theta$, such that $R(Z_{\theta})=AZ_{\theta}B$ (where $R$ is applied entrywise) and $A^{k}=B^{k}=I$. The antipodal action ($k=2$) is particularly simple; in this case each entry of $Z_{\theta}$ is odd, so $R(Z_{\theta})=-Z_{\theta}$. See [39, §3] for a slightly modified construction that chooses the monomial coefficients continuously.
@@
The $\theta$-spheres are related to a similar family of $C^{*}$-algebras called quantum (or noncommutative) tori ([20, §12.2] or [48, §1]). If $\theta$ is an antisymmetric $n\times n$ real matrix, the $C^{*}$-algebra $A^{n}_{\theta}$, or $C(\mathbb{T}^{n}_{\theta})$, is defined by generators $u_{1},\ldots,u_{n}$ and relations
@@
All $\theta$-spheres and noncommutative tori are known to be Rieffel deformations in the sense of [50, §10.2]. While it is not clear if the $\theta$-spheres or noncommutative tori form a continuous field over the set of antisymmetric matrices, [50, Theorem 8.13] implies that the field is separately continuous in each entry. In particular, the distinction between those two cases does not matter for $3$-spheres or $2$-tori. As we will discuss continuity of fields in more detail, we recall some basic definitions.
@@
$C(X)$-algebras are one way to formalize the notion of a field of $C^{*}$-algebras [47, Definition 1.1 and Propositions 1.2 and 1.3], while [37, Theorem 2.3] equates the notion to that of an upper semicontinuous $C^{*}$-algebra bundle, generalizing [18, Definition VII.8.2].
@@ 3 Local-Triviality Dimension and Fields
The local-triviality dimension $\mathrm{dim}^{{\mathbb{G}}}_{\mathrm{LT}}(A)$ is the infinimum of the set of $n\in\mathbb{N}$ such that there exist ${\mathbb{G}}$-equivariant *-homomorphisms $\rho_{0},\ldots,\rho_{n}:C_{0}((0,1])\otimes C({\mathbb{G}})\to A$ with $\sum\limits_{j=0}^{n}\rho_{j}(t\otimes 1)=1$. $\blacklozenge$
@@
The weak local-triviality dimension $\mathrm{dim}^{{\mathbb{G}}}_{\mathrm{WLT}}(A)$ is the infinimum of the set of $n\in\mathbb{N}$ such that there exist ${\mathbb{G}}$-equivariant *-homomorphisms $\rho_{0},\ldots,\rho_{n}:C_{0}((0,1])\otimes C({\mathbb{G}})\to A$ with $\sum\limits_{j=0}^{n}\rho_{j}(t\otimes 1)$ invertible. $\blacklozenge$
@@
The strong local-triviality dimension is the infimum of the set of $n\in\mathbb{N}$ such that there exists a ${\mathbb{G}}$-equivariant unital *-homomorphism $\rho:C(E_{n}^{\Delta}{\mathbb{G}})\to A$. $\blacklozenge$
@@
Multiple characterizations of freeness are also found in [4, Theorem 0.4]. Moreover, finiteness of any one of the local-triviality dimensions implies freeness of the coaction by [19, Theorem 3.8]. We will primarily be interested in the case that $A$ is noncommutative, but ${\mathbb{G}}$ is an ordinary group (in particular, a compact abelian group). In this case, there is an equivalent condition to freeness found in [49, Definition 1.6] or [44, Definition 7.1.4]. If ${\mathbb{G}}$ is compact abelian and $\alpha$ is an action of ${\mathbb{G}}$ on $A$, then for each character $\gamma\in\widehat{{\mathbb{G}}}$, there is an associated eigenspace (or spectral subspace)
@@
Immediate, as there exist ${\mathbb{G}}$-equivariant morphisms $A\to A_{x}$ for all $x\in X$. $\blacksquare$
@@
is thus a self-map of the sphere that has no fixed points and sends no point to its antipode, and no such map exists by [58, Corollary 1.24]. $\blacklozenge$
@@
Equip ${\mathbb{P}}^{1}$ with its complex analytic structure or, equivalently for the purpose of manipulating holomorphic bundles, its complex projective algebraic structure (see [53, §12]). The line bundle
@@
Having to work with such trivializations extensively, we need some language to express the extent to which they exist. To that end, the following notions specialize the (numerical) ${\mathbb{G}}$-index ([62, Definition 3.13], [30, Definition 6.2.3]) of a ${\mathbb{G}}$-action on a space $X$, which is typically compact. That invariant can be recast via [13, Theorem 3.4] (where the term is co-index instead) as the smallest $n$ for which $X/{\mathbb{G}}$ admits an open cover by $U_{0},\ldots,U_{n}$, above which the action induces trivial principal ${\mathbb{G}}$-bundles.
@@
For a ${\mathbb{Z}}/k$-action on a matrix bundle ${\mathcal{M}}$ over $X$, the corresponding index $\operatorname{ix}({\mathcal{M}},\triangleright)=\operatorname{ix}(\triangleright)$ is the smallest $m$ for which there is an open cover $X=\bigcup_{i=0}^{m}U_{i}$ with ${\mathcal{M}}$ trivial, $\triangleright$ inner over each $U_{i}$ induced by some unitary $s_{i}\in\Gamma({\mathcal{M}}|_{U_{i}})$, $s_{i}^{k}=1$, with all eigenspaces of $s_{i}$ again trivial as bundles over $U_{i}$. $\blacklozenge$
@@
Given a ${\mathbb{Z}}/k$-action $\triangleright$ on a matrix bundle ${\mathcal{M}}\to X$ over a compact Hausdorff space, we have
$\operatorname{\mathrm{qt}}-1\,\,\leq\,\,\mathrm{dim}^{{\mathbb{Z}}/k}_{\mathrm{WLT}}(\Gamma({\mathcal{M}}))\,\,\leq\,\,\operatorname{\mathrm{qt}}\cdot\left(\operatorname{ix}({\mathcal{M}},\triangleright)+1\right)-1.$ (3-4)
@@
First consider when $X$ is a singleton, so $\operatorname{ix}({\mathcal{M}},\triangleright)=0$ and ${\mathcal{M}}$ is a matrix algebra $M_{n}$. The result [12, Theorem 4.1] effectively proves that $\mathrm{dim}^{{\mathbb{Z}}/k}_{\mathrm{WLT}}(M_{n})=\operatorname{\mathrm{qt}}-1$. Although that discussion focuses on gauge actions on matrix algebras, the arguments therein still work for arbitrary conjugation actions by unitaries $v\in M_{n}$, $v^{k}=1$.
Applying this result to general bundles shows that each individual fiber $\Gamma({\mathcal{M}})_{x}\cong M_{n_{x}}$ satisfies $\mathrm{dim}^{{\mathbb{Z}}/k}_{\mathrm{WLT}}(\Gamma({\mathcal{M}})_{x})=\operatorname{\mathrm{qt}}_{x}-1$. Because there is an equivariant map
$\Gamma({\mathcal{M}})\xrightarrow{\quad}\Gamma({\mathcal{M}})_{x}$
for each $x\in X$, we obtain that
$\mathrm{dim}^{{\mathbb{Z}}/k}_{\mathrm{WLT}}(\Gamma({\mathcal{M}}))\,\,\geq\,\,\sup_{x\in X}\,\mathrm{dim}^{{\mathbb{Z}}/k}_{\mathrm{WLT}}(\Gamma({\mathcal{M}})_{x})\,\,=\,\,\sup_{x\in X}\,\operatorname{\mathrm{qt}}_{x}-1\,\,=\,\,\operatorname{\mathrm{qt}}-1,$
which establishes the lower bound on the weak dimension.
For the upper bound on the weak dimension, we may assume $\operatorname{\mathrm{qt}}$ is finite, as otherwise the result is trivial. Also, define $\operatorname{ix}:=\operatorname{ix}({\mathcal{M}},\triangleright)$ for simplicity, and consider an open cover $X=\bigcup_{i=0}^{\operatorname{ix}}U_{i}$ witnessing Definition 3.12. That is, the open cover trivializes the matrix bundle ${\mathcal{M}}$ and renders the action inner, by unitary sections $v_{i}\in\Gamma({\mathcal{M}}|_{U_{i}})$ with $v_{i}^{k}=1$, and with trivial $v_{i}$-eigenbundles
${\mathcal{E}}_{i,j}:=\ker(\zeta^{j}-v_{i}),\quad\quad 0\leq i\leq\operatorname{ix},\quad j\in{\mathbb{Z}}/k.$
If $i$ is fixed, the various ${\mathcal{E}}_{i,j}$ are uniquely determined up to cycling through the $j$ indices.
Because $\operatorname{\mathrm{qt}}$ is assumed finite, each ${\mathcal{E}}_{i,j}$ has nonzero vector space dimension at each $x\in X$. If $i$ is fixed and $x\in U_{i}$ is fixed, consider the smallest such eigenspace $W_{0}:=({\mathcal{E}}_{i,j_{0}})_{x}$. It is possible to define a partial isometry on $M_{n_{x}}$ that maps $W_{0}$ isometrically onto a subspace $W_{1}\subseteq({\mathcal{E}}_{i,j_{0}+1})_{x}$ (which is guaranteed to be orthogonal to $W_{0}$ because the eigenspaces are orthogonal), maps $W_{1}$ isometrically onto some $W_{2}\subseteq({\mathcal{E}}_{i,j_{0}+2})_{x}$, and so on, eventually returning to $W_{k}=W_{0}=({\mathcal{E}}_{i,j_{0}})_{x}$. Further, map every element of $\left(\bigoplus_{m=0}^{k-1}W_{m}\right)^{\perp}$ to 00. This procedure produces a partial isometry $\alpha$ in the generating spectral subspace of the action of ${\mathbb{Z}}/k$ on this fiber. In fact, $\alpha$ is normal, as the initial space and final space are both equal to the direct sum $\bigoplus_{m=0}^{k-1}W_{m}$, and it also follows from the construction that $\alpha^{k}$ is equal to the projection onto that same direct sum. The subspaces $W_{m}$ for $m>0$ may be proper subspaces of the corresponding eigenspaces, but because $\operatorname{\mathrm{qt}}$ is defined as the (largest) quotient spread, we need at most $\operatorname{\mathrm{qt}}$ such operators $\alpha_{1},\ldots,\alpha_{\operatorname{\mathrm{qt}}}$ to guarantee full coverage of each eigenspace. That is, $\sum\limits_{s=1}^{\operatorname{\mathrm{qt}}}\alpha_{s}^{*}\alpha_{s}$ is an invertible element of $M_{n_{x}}$.
Now, consider that since the eigenbundles ${\mathcal{E}}_{i,j}$ are trivial on $U_{i}$, we may apply this construction on each $U_{i}$, rather than at an individual point. This produces normal partial isometries $\alpha_{i,s}\in\Gamma({\mathcal{M}}|_{U_{i}})$, for $0\leq i\leq\operatorname{ix}$ and $1\leq s\leq\operatorname{\mathrm{qt}}$, that map
${\mathcal{E}}_{i,j}\xrightarrow{\quad\alpha_{i,s}\quad}{\mathcal{E}}_{i,j+1},\quad j\in{\mathbb{Z}}/k$
and satisfy the conditions that each $\alpha_{i,s}^{k}$ is a projection and $\sum\limits_{s=1}^{\operatorname{\mathrm{qt}}}\alpha_{i,s}^{*}\alpha_{i,s}$ is invertible on each $U_{i}$. As before, we have that the $\alpha_{i,s}$ are in the generating spectral subspace of the ${\mathbb{Z}}/k$ action by design. However, each $\alpha_{i,s}$ is only a section on $U_{i}$, not the full space $X$.
Finally, we may fix a refinement $(V_{i})_{i=0}^{\operatorname{ix}}$ of the cover $(U_{i})_{i=0}^{\operatorname{ix}}$, with
@@
A partition of unity argument (see [34, Theorem 36.1]) allows us to rescale the operators $\alpha_{i,s}\in\Gamma({\mathcal{M}}|_{U_{i}})$ into global sections $\phi_{i}^{1/2}\alpha_{i,s}\in\Gamma(M)$, where the $\phi_{i}$ are nonnegative scalar-valued continuous functions on $X$, supported in $U_{i}$, with $\phi_{i}|_{V_{i}}\equiv 1$ and $\sum\limits_{i=0}^{\operatorname{ix}}\phi_{i}\equiv 1$. Note that each $\phi_{i}$ is a central element of $\Gamma({\mathcal{M}})$ and is fixed by the action of ${\mathbb{Z}}/k$, so multiplying by $\phi_{i}^{1/2}$ has not removed any of the key properties of $\alpha_{i,s}$. Because $1\leq s\leq\operatorname{\mathrm{qt}}$ and $0\leq i\leq\operatorname{ix}$, we have produced a list of $\operatorname{\mathrm{qt}}\cdot(\operatorname{ix}+1)$ normal operators $\alpha_{i,s}$ in the generating spectral subspace of the ${\mathbb{Z}}/k$ action such that $\sum\limits_{i,s}\alpha_{i,s}^{*}\alpha_{i,s}$ is (globally) invertible. By Proposition 3.5, we have that
$\mathrm{dim}^{{\mathbb{Z}}/k}_{\mathrm{WLT}}(\Gamma({\mathcal{M}}))\,\,\leq\,\,\operatorname{\mathrm{qt}}\cdot\left(\operatorname{ix}+1\right)-1.$
$\blacksquare$
@@
For a ${\mathbb{Z}}/k$-action $\triangleright$ on a matrix bundle ${\mathcal{M}}\to X$ over a compact Hausdorff space, the following conditions are equivalent.
@@
The implication a $\Longrightarrow$ b follows from the upper bound 3-4 on $\mathrm{dim}^{{\mathbb{Z}}/k}_{\mathrm{WLT}}(\Gamma({\mathcal{M}}))$, where we note that $\operatorname{ix}({\mathcal{M}},\triangleright)$ is always finite, and the full spectrum condition of (a) guarantees that the definition of $\operatorname{\mathrm{qt}}$ does not involve dividing by zero. Moreover, b $\Longrightarrow$ c follows from the general result [19, Theorem 3.8]. As for c $\Longrightarrow$ a, freeness means by [33, Theorem 8.1.7] that the ${\mathbb{Z}}/k$-grading on $A:=\Gamma({\mathcal{M}})$ induced by the action is strong:
@@
then the orthogonal projection onto the latter (non-zero) space annihilates $M_{n,\zeta}M_{n,\zeta^{-1}}$. $\blacksquare$
@@
For a ${\mathbb{Z}}/k$-action $\triangleright$ on a matrix bundle ${\mathcal{M}}\to X$ over a compact Hausdorff space, the following conditions are equivalent.
@@
    $\mathrm{dim}^{{\mathbb{Z}}/k}_{\mathrm{LT}}(\Gamma({\mathcal{M}}))\leq\operatorname{ix}({\mathcal{M}},\triangleright)$.
@@
    $\mathrm{dim}^{{\mathbb{Z}}/k}_{\mathrm{LT}}(\Gamma({\mathcal{M}}))<\infty$.
@@
The implication a $\Longrightarrow$ b is formal, and b $\Longrightarrow$ c follows from the proof of [12, Theorem 4.4], which provides the $X=\{*\}$ version. We thus turn to c $\Longrightarrow$ a.
As in the proof of Theorem 3.13, consider an open cover $X=\bigcup_{i=0}^{\operatorname{ix}}U_{i}$ trivializing ${\mathcal{M}}$, generators $v_{i}$ for the ${\mathbb{Z}}/k$-action restricted to $U_{i}$, and the eigenspaces ${\mathcal{E}}_{i,j}$, $j\in{\mathbb{Z}}/k$ of $v_{i}$. We can construct operators $\alpha_{i,s}$ as in the earlier proof, with the only difference being that now we can select a single, unitary $\alpha_{i}\in\Gamma({\mathcal{M}}|_{U_{i}})$ for each $i$. The same partition of unity argument will produce a list of $\operatorname{ix}({\mathcal{M}},\triangleright)+1$ operators in the vein of Proposition 3.5, demonstrating that $\mathrm{dim}^{{\mathbb{Z}}/k}_{\mathrm{LT}}(\Gamma({\mathcal{M}}))\leq\operatorname{ix}({\mathcal{M}},\triangleright)$. $\blacksquare$
@@
$\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{WLT}}(A)=\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{LT}}(A)=1\quad\text{and}\quad\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{SLT}}(A)=\infty$
@@
Because (by oddness) each eigenspace of $a_{i}$ is orthogonal to each eigenspace of $u$, 3-5 coincides with the canonical surjection ${\mathbb{S}}^{2}\to{\mathbb{R}}{\mathbb{P}}^{2}$ nowhere. A continuous lift ${\mathbb{S}}^{2}\to{\mathbb{S}}^{2}$ thereof (which does exist by [34, Lemma 79.1], as ${\mathbb{S}}^{2}$ is simply-connected) will thus have no coincidences with either the identity or the antipodal map, once more violating [58, Corollary 1.24]. $\blacksquare$
@@
by projectivity, as in the proof of [12, Proposition 5.4], via [45, Proposition 2.10]. Define $a:=\sum\limits_{i=0}^{n}\overline{\rho}_{i}(t\otimes 1)$. Note that $\pi_{x}(a)$ is invertible, so we may choose an element $b\in A$ such that $\pi_{x}(b)=(\pi_{x}(a))^{-1}$. That is, both $\|ab-1\|_{x}=0$ and $\|ba-1\|_{x}=0$. By [47, Proposition 1.2], for each $c\in A$ the mapping